A pole vaulter approaches the takeoff point at a speed of 8.81m/a. Assuming that only this speed determines the height to which they can rise, find the maximum height which the vaulter can clear the bar.


1
Expert's answer
2021-04-29T10:43:49-0400

Use conservation of energy to find the height:


"KE=PE,\\\space\\
\frac 12mv^2=mgh,\\\space\\
h=\frac{v^2}{2g}=3.96\text{ m}."